OpenStudy (anonymous):

bob has five times as much money as john and together they have $60.00

OpenStudy (tkhunny):

Okay, what's your plan? Give Bob's money a name.

TheSmartOne (thesmartone):

Ok lets say Bob's money is B and John's money is J.

TheSmartOne (thesmartone):

Bob has 5 times as much as John. So for example if John has $2 then Bob will have $10 So in that case we can make a relation between them whih is: B=5*J

TheSmartOne (thesmartone):

Abd B+J=60

TheSmartOne (thesmartone):

So we get 2 equations B+J=60 B=5*J

TheSmartOne (thesmartone):

You can solve them by substituting.

TheSmartOne (thesmartone):

Can you solve from there? @minibarbie

OpenStudy (tkhunny):

Gaaa!!!! Now, we'll never know what @minibarbie would have named Bob's money.